OpenStudy (britbrat1997):

Solve the inequality and write your answer in interval notation. 0.1x+0.8≤13.8+0.7x

OpenStudy (phi):

you could add -0.1x to both sides. what do you get ?

jhonyy9 (jhonyy9):

multiplie both sides by 10 for you c

jhonyy9 (jhonyy9):

can eliminate the decimals

jhonyy9 (jhonyy9):

with integers you can calculi more easy

OpenStudy (jdoe0001):

ditto so... how would "you" solve ... say 0.1x+0.8 = 13.8+0.7x ?

OpenStudy (britbrat1997):

0.1x+0.8≤13.8+0.7x 0.8<=13.8+.6x -13<=0.6x

OpenStudy (britbrat1997):

12.67=x?

OpenStudy (britbrat1997):

how do i write it in interval notation?

OpenStudy (jdoe0001):

12.67... actually close but yes, you'd do the same more or less, as you'd do in an equation

OpenStudy (jdoe0001):

\(\bf 0.1x+0.8\le13.8+0.7x\implies 0.8\le 13.8+0.6x \\ \quad \\ -13\le 0.6x\implies \cfrac{-13}{0.6}\le x\)

OpenStudy (jdoe0001):

the only difference is that, in inequalities, when you multiply or divide or exponentialize by a negative number, you have to "flip" the inequality sign notice above, we divided by 0.6, BUT 0.6 is positive, so the sign remains \(\large \le\)

OpenStudy (jdoe0001):

to use decimals -13/0.6 is -21.6666 so, the inequaliy is saying "x" is a value that greater than or equals to -21.6666 recall that on the negative side, the closer to the 0, the bigger the number so -19 is bigger than -21, because -19 is closer to the 0 so, interval notation wise